OpenStudy (anonymous):

What is the Greatest common factor of 24 and 32, please include a factors list for each.

OpenStudy (alexandervonhumboldt2):

24/2=12 32/2=16 12/2=6 16/2=8 6/2=3 8/2=4 3/3=1 4/2=2 2/2=1 2.2.2.2.2 2,2,2,3 the GCF is 2*2*2=8